# Toward Efficient Spiking Transformers: Synapse Pruning Meets Synergistic Learning-Based Compensation

### v2

@@ Abstract
Figure 1: Comparison of model parameters and classification accuracy between our lightweight models and existing ST-based models on the CIFAR10-DVS dataset.
Introduction
@@
The Transformer architecture (Vaswani et al. 2017) has emerged as a foundational backbone for a wide array of large language models (Achiam et al. 2023; Devlin et al. 2019; Touvron et al. 2023), owing to its strengths in modeling long-range dependencies, adaptability across multiple modalities, and high parallelism. Recently, researchers have extended Transformer architectures to spiking neural networks (SNNs), giving rise to spiking Transformer (ST)-based models (Zhou et al. 2023b; Yao et al. 2023a). By making targeted adaptations to the dynamics of spiking neurons, these models have demonstrated significant improvements across a variety of complex tasks (Yao et al. 2024; Zhou et al. 2024a). However, the ST is inherently a parameter-intensive architecture, where performance often correlates with the model size (Dosovitskiy et al. 2021), resulting in substantial computational and efficiency costs. Achieving parameter-efficient adaptations and low computational cost thus remains a critical challenges in this field.
Figure 2: Overview of the proposed lightweight ST-based models and the associated pruning strategies.
@@
To enhance the efficiency of spatio-temporal (ST) blocks, numerous methods have been proposed in prior work. Among these, structure-oriented optimization represents a key direction, focusing on architectural modifications such as spatio-temporal pruning (Zhou et al. 2024b), meta-architecture redesign (Yao et al. 2024), and the integration of lightweight functional modules (Zhou et al. 2024a). These techniques aim to improve representational efficiency within ST blocks. Additionally, engineering-oriented approaches—such as quantization-aware training (Qiu et al. 2025) and specialized neuromorphic processors (Kim et al. 2025)—address practical concerns like energy consumption and inference latency, facilitating the deployment of ST-based models on resource-constrained edge devices. Despite their effectiveness, these approaches are often tightly coupled with specific model architectures or constrained by task-specific assumptions, limiting their generalizability (refer to Related Work in the Appendix for additional details). In contrast, our goal is to develop a universal, model-agnostic compression strategy for ST blocks without compromising performance.
@@
SNNs are composed of biologically inspired spiking neurons that emulate the spatio-temporal dynamics observed in human brain activity. Drawing inspiration from these biological mechanisms, researchers have proposed a range of enhanced spiking neuron models (Qiu et al. 2025; Huang et al. 2024; Wang et al. 2023; Fang et al. 2021). Notably, models such as IE-LIF (Qiu et al. 2025), multi-threshold (Huang et al. 2024), and dynamic-threshold neurons (Wang et al. 2023) have been integrated into ST blocks to improve the efficiency of spiking units (see the Related Work in the Appendix for details). In this work, we aim to proposing an enhanced spiking neuron model that is efficient and simple to implement and train. We hope that the enhanced spiking neuron model can improve the efficiency of ST blocks, leading to more lightweight ST-based models.
@@
Our method address two key advantages. First, the pruning pipeline allows flexible control over model size by adjusting the desired sparsity level. Second, the proposed sLIF neuron is plug-and-play, enabling seamless integration into existing ST-based models. As shown in Fig. 1, our approach achieves higher accuracy with fewer parameters compared to existing ST-based models on the neuromorphic CIFAR10-DVS dataset.
The contributions of this work are as follows:
@@
  • An enhanced sLIF neuron model is proposed to effecitvely compensate for the information loss based on the synergistic learing between synaptic and intrinsic plasticity.
@@
Methods
@@
Efficient Compression for Spiking Transformer
@@
Here, the SSA and MLP represent the spiking self-attention module and linear layers module, respectively. In a ST-based model, the spatio-temporal patch embeddings $x_{0}\in\mathbb{R}^{T\times N\times d}$ are progressively refined through $L$ stacked $\mathrm{ST}$ blocks. The resulting high-level representations are subsequently fed into a task-specific prediction head to produce the final output.
@@
$\mathbf{L_{1}}$ Norm-based Parameter Sparsification
Unstructured pruning offers a straightforward and flexible approach to model compression, with high implementation adaptability and demonstrated effectiveness in previous research. In our method, referred to as $\mathbf{L_{1}P}$, we prune by zeroing out a specified proportion of elements with the lowest $\mathrm{L_{1}}$-norm values within each target weight matrix. This strategy effectively reduces the number of model parameters while preserving the overall functional integrity of the network.
@@
Dimension Significance-based Pruning
@@
Synergistic Learning-based Compensation
@@
Various spiking neuron models have been proposed to mimic the spatio-temporal dynamics of biological neurons. Among them, the LIF model is widely used in SNNs due to its balance between biological fidelity and computational efficiency. With surrogate gradient methods like spatio-temporal backpropagation (STBP), SNNs have shown notable performance in learning complex patterns. However, most studies focus on synaptic plasticity, neglecting the role of intrinsic neuronal parameters in modulating excitability. To bridge this gap, we propose the sLIF model, which integrates synaptic and intrinsic plasticity (IP) into a unified learning framework.
The dynamics of the proposed sLIF neuron are governed by the following differential equation:
@@
where $u_{t}$ denotes the membrane potential at time $t$, $u_{\mathrm{rest}}$ is the resting potential, and $I_{t}$ is the total synaptic input current received by the neuron. The generation of output spikes follows a thresholding mechanism:
@@
where $\mathrm{H}(\cdot)$ denotes the Heaviside step function, and $u_{\mathrm{th}}$ is the firing threshold. The synaptic input $I_{t}$ is computed as the weighted summation of incoming spikes:
$I_{t}=\sum_{i}I_{\mathrm{syn}}^{i,t}=\sum_{i}w^{i}o^{i,t},$ (13)
@@
Distinct from prior models, the sLIF neuron treats intrinsic parameters, the membrane time constant $\tau$ and the firing threshold $u_{\mathrm{th}}$, as learnable parameters, and optimizes them alongside the synaptic weights $w$. As illustrated in Fig. 3, the membrane time constant determines the rate at which historical information is forgotten, while the firing threshold directly influences the neuron’s firing rate. The synergistic learning between synaptic and intrinsic parameters enhances neuronal heterogeneity, thereby improving the representational capacity of the SNN models.

### v4

@@ Abstract
publicationid: pubid: This work has been submitted to the IEEE for possible publication.
Copyright may be transferred without notice, after which this version may no longer be accessible.
@@
Figure 1: Comparison of model parameters and classification accuracy between our lightweight models (denoted as ‘ST[ours]’) and existing ST-based models on the CIFAR10-DVS dataset.
I Introduction
@@
ThE Transformer architecture [vaswani2017attention] has emerged as a foundational backbone for a wide array of large language models [achiam2023gpt, touvron2023llama], owing to its strengths in modeling long-range dependencies, adaptability across multiple modalities, and high parallelism. Recently, researchers have extended Transformer architectures to spiking neural networks (SNNs), giving rise to spiking Transformer (ST)-based models [zhou2023spikformer, yao2023spike]. By making targeted adaptations to the dynamics of spiking neurons, these models have demonstrated significant improvements across a variety of complex tasks [yao2024spikedriven, zhou2024qkformer, 10772601]. However, the ST is inherently a parameter-intensive architecture, where performance often correlates with the model size [dosovitskiy2021an, pnas2528527122], resulting in substantial computational and efficiency costs. Achieving parameter-efficient adaptations and low computational cost thus remains a critical challenge in this field.
@@
To enhance the efficiency of ST blocks, numerous methods have been proposed in prior work. Among these, structure-oriented optimization represents a key direction, focusing on architectural modifications such as spatio-temporal pruning [zhou2024spatial], meta-architecture redesign [yao2024spikedriven], and the integration of lightweight functional modules [zhou2024qkformer]. These techniques aim to improve representational efficiency within ST blocks. Additionally, engineering-oriented approaches—such as quantization-aware training [qiu2025quantized] and specialized neuromorphic processors [kim2025c]—address practical concerns like energy consumption and inference latency, facilitating the deployment of ST-based models on resource-constrained edge devices. Despite their effectiveness, these approaches are often tightly coupled with specific model architectures or constrained by task-specific assumptions, limiting their generalizability. In contrast, our goal is to develop a universal, model-agnostic compression strategy for ST blocks without compromising performance.
@@
SNNs are composed of biologically inspired spiking neurons that emulate the spatio-temporal dynamics observed in human brain activity. Drawing inspiration from these biological mechanisms, researchers have proposed a range of enhanced spiking neuron models [qiu2025quantized, huang2024towards, wang2023complex, fang2021incorporating]. Notably, models such as IE-LIF [qiu2025quantized], multi-threshold [huang2024towards], and dynamic-threshold neurons [wang2023complex] have been integrated into ST blocks to improve the efficiency of spiking units. In this work, we aim to propose an enhanced spiking neuron model that is efficient and simple to implement and train. We hope that the enhanced spiking neuron model can improve the efficiency of ST blocks, leading to more lightweight ST-based models.
@@
Our method addresses two key advantages. First, the pruning pipeline allows flexible control over model size by adjusting the desired sparsity level. Second, the proposed sLIF neuron is plug-and-play, enabling seamless integration into existing ST-based models. As shown in Fig. 1, our approach achieves higher accuracy with fewer parameters compared to existing ST-based models on the neuromorphic CIFAR10-DVS dataset.
The main contributions and highlights of this study can be summarized as follows.
@@
  • An enhanced sLIF neuron model is proposed to effectively compensate for the information loss based on the synergistic learning between synaptic and intrinsic plasticity.
@@
The remainder of this article is organized as follows. Section II reviews prior studies on efficient spiking transformers and bio-inspired spiking neuron models. Section III introduces the proposed lightweight strategy for spiking transformers and details its compensation mechanism based on synergistic learning. Section IV describes the experimental framework, including the setup, results, and corresponding analyses. Finally, Section V summarizes the key findings and concludes the paper.
II Related Work
In this section, we briefly review recent works on efficient spiking transformers and bio-inspired spiking neuron models that are closely related to our study.
II-A Efficient Spiking Transformers
ST-based models have garnered significant attention due to their potential for biologically plausible computation and energy-efficient processing of spatio-temporal data [zhou2023spikformer, yao2023spike]. However, the high computational cost and memory overhead inherent in their complex dynamics necessitate efficiency-oriented enhancements. Prior work can be broadly categorized into structure-oriented and engineering-oriented approaches.
Structure-oriented optimization: A primary direction for improving efficiency in ST blocks lies in architectural redesign. Inspired by model pruning strategies, researchers have explored eliminating redundant spatial and temporal components in attention mechanisms and feedforward layers, thereby reducing computational burden without substantially degrading accuracy [zhou2024spatial]. Another line of work proposes meta-architecture redesigns [yao2024spikedriven, wang2025spiking], which tailor model structures to the unique characteristics of spiking data. Furthermore, the integration of lightweight functional modules—such as linearized QK-value embedding and simplified gating mechanisms—have shown promise in improving representational efficiency with minimal performance trade-offs [zhou2024qkformer].
Engineering-oriented solutions: In parallel, engineering-based methods addressing practical deployment issues have been proposed. Weight quantization, which reduces numerical precision, is a widely used technique to enable model deployment on edge devices. Researchers have incorporated quantization-aware training into ST-based models to lower the bit-width of weights and activations, thereby minimizing memory footprint and enabling deployment on low-power hardware [qiu2025quantized]. Similarly, advances in specialized neuromorphic hardware—such as event-driven computing cores and custom-designed crossbar arrays—have significantly reduced latency and energy consumption for ST-based models [kim2025c]. Despite their success, these methods tend to be model-dependent or require non-trivial system-level integration. This lack of universality limits their adaptability to diverse architectures and downstream tasks.
II-B Bio-Inspired Spiking Neuron Models
The spiking neuron, as the fundamental computational unit, plays a critical role in determining the overall performance of SNNs [zheng2024temporal]. Among various neuron models, the LIF model is one of the most widely adopted due to its simplicity and computational efficiency. By training LIF neurons with synaptic plasticity mechanisms [wu2018spatio], SNN models have achieved competitive performance on various tasks [sun2024reliable, 10318216]. Despite these advantages, the limited representational capacity of LIF in capturing complex spatio-temporal dynamics has motivated the development of more biologically inspired neuron models.
In spiking neurons, intrinsic parameters (i.e., the membrane time constant, firing threshold, and resting potential) play a crucial role in shaping internal dynamics, including input integration, spike generation, and the refractory period [9449951, klinshov2024extending]. Based on the biological mechanisms underlying the firing threshold, numerous enhanced variants of the LIF model have been proposed, focusing on the threshold adaptation, dynamic thresholds, and subthreshold dynamics [sun2023synapse, huang2024towards]. Compared to the standard LIF neuron, these models offer a more biologically plausible depiction of threshold behavior, thereby enhancing the neuronal ability to robustly represent spatial patterns. Furthermore, the membrane time constant governs the temporal integration of incoming signals [pazderka2024role]. By introducing heterogeneity in this parameter, recent studies have significantly improved the capacity of SNNs to capture temporal features across multiple time scales [fang2021incorporating, 10857949].
A key challenge in enhancing LIF neuron models lies in the effective integration of the synergistic interactions between synaptic and intrinsic parameters. Prior studies have typically treated these parameters as independent components or optimized them using information-theoretic approaches, often overlooking their interdependence [li2013synergies, 10235316]. This not only limits the model expressiveness but also complicates the training process. Therefore, two critical questions remain open: how to incorporate synaptic and intrinsic plasticity in a unified and efficient framework, and how to enable synergistic learning between these two forms of plasticity.
Figure 2: Overview of the proposed lightweight ST-based models and the corresponding pruning strategies. In the original ST-based encoder, input embeddings are sequentially processed by the SSA and MLP modules. The primary parameter overhead resides in matrices $\mathbf{U}{\mathrm{q}}$, $\mathbf{U}{\mathrm{k}}$, $\mathbf{U}{\mathrm{v}}$, and $\mathbf{M}{0}$ of the SSA module, as well as $\mathbf{M}{1}$ and $\mathbf{M}{2}$ of the MLP module. To address this, two lightweight strategies are introduced: L1P, which yields sparse matrices, and DSP, which produces low-rank matrices.
III Methods
@@
III-A Efficient Compression for Spiking Transformer
@@
Here, the SSA and MLP represent the spiking self-attention module and linear layers module, respectively. In an ST-based model, the spatio-temporal patch embeddings $x_{0}\in\mathbb{R}^{T\times N\times d}$ are progressively refined through $L$ stacked $\mathrm{ST}$ blocks. The resulting high-level representations are subsequently fed into a task-specific prediction head to produce the final output.
@@
III-A1 $\mathbf{L_{1}}$ Norm-based Parameter Sparsification
Unstructured pruning offers a straightforward and flexible approach to model compression, with high implementation adaptability and demonstrated effectiveness in previous research. In our method, referred to as $\mathbf{L_{1}P}$, we prune by zeroing out a specified proportion of elements with the smallest $\mathrm{L_{1}}$-norm values within each target weight matrix. This strategy effectively reduces the number of model parameters while preserving the overall functional integrity of the network.
@@ III-A1 $\mathbf{L_{1}}$ Norm-based Parameter Sparsification
III-A2 Dimension Significance-based Pruning
@@
III-B Synergistic Learning-based Compensation
@@
Various spiking neuron models have been proposed to mimic the spatio-temporal dynamics of biological neurons. Among them, the LIF model is widely used in SNNs due to its balance between biological fidelity and computational efficiency. With surrogate gradient methods like spatio-temporal backpropagation (STBP), SNNs have shown notable performance in learning complex patterns. However, most studies focus on synaptic plasticity, neglecting the role of intrinsic neuronal parameters in modulating excitability. As illustrated in Fig. 3, the membrane time constant determines the rate at which historical information is forgotten, while the firing threshold directly influences the neuronal firing rate. The synergistic learning between synaptic and intrinsic parameters can enhance neuronal heterogeneity, thereby improving the representational capacity of the SNN models.
To bridge this gap, we propose the sLIF model, which integrates synaptic and intrinsic plasticity (IP) into a unified learning framework. The dynamics of the proposed sLIF neuron are governed by the following differential equation:
@@
where $u_{t}$ denotes the membrane potential at time $t$, $u_{\mathrm{rest}}$ is the resting potential, and $I_{t}$ is the total synaptic input current received by the neuron. In practice, we discretize Eq. (11) using the Euler method with a unit time step for computational implementation:
$u_{t}=u_{t-1}+\frac{-(u_{t-1}-u_{\mathrm{rest}})+I_{t}}{\tau}.$ (12)
The generation of output spikes follows a thresholding mechanism:
@@
where $\mathrm{H}(\cdot)$ denotes the Heaviside step function, and $u_{\mathrm{th}}$ is the firing threshold. Following a spike, the membrane potential is reset using either a soft-reset mechanism:
$u_{t}=u_{t}-o_{t}u_{\mathrm{th}},$ (14)
or a hard-reset mechanism:
$u_{t}=u_{t}(1-o_{t}).$ (15)
The synaptic input $I_{t}$ is computed as the weighted summation of incoming spikes:
$I_{t}=\sum_{i}w^{i}o^{i}_{t},$ (16)
@@
Distinct from prior models, the sLIF neuron treats intrinsic parameters, the membrane time constant $\tau$ and the firing threshold $u_{\mathrm{th}}$, as learnable parameters, and optimizes them alongside the synaptic weights $w$. According to the chain rule, the derivatives of the loss function $Loss$ with respect to $o^{i}_{t,n}$ and $u^{i}_{t,n}$ can be mathematically described as follows:
$\frac{\partial Loss}{\partial o^{i}_{t,n}}=\frac{\partial Loss}{\partial o^{i}_{t+1,n}}\frac{\partial o^{i}_{t+1,n}}{\partial o^{i}_{t,n}}+\sum_{j=1}^{l(n+1)}\frac{\partial Loss}{\partial o^{j}_{t,n+1}}\frac{\partial o^{j}_{t,n+1}}{\partial o^{i}_{t,n}},$ (17)
$\frac{\partial Loss}{\partial u^{i}_{t,n}}=\frac{\partial Loss}{\partial o^{i}_{t,n}}\frac{\partial o^{i}_{t,n}}{\partial u^{i}_{t,n}}+\frac{\partial Loss}{\partial o^{i}_{t+1,n}}\frac{\partial o^{i}_{t+1,n}}{\partial u^{i}_{t,n}}.$ (18)
Here, $o^{i}_{t,n}$ and $u^{i}_{t,n}$ represent the output spike and membrane potential of the $i$-th neuron in the $n$-th layer at time $t$. $l(n+1)$ denotes the number of neurons in the $(n+1)$-th layer. Based on equations (17) and (18), we finally obtain the derivatives with respect to the synaptic weight, firing threshold and membrane time constant:
$\frac{\partial Loss}{\partial w^{i}_{n}}=\sum_{t=1}^{T}\frac{\partial Loss}{\partial u^{i}_{t,n}}\frac{\partial u^{i}_{t,n}}{\partial w^{i}_{n}},$ (19)
$\frac{\partial Loss}{\partial u_{\text{th},n}}=\sum_{t=1}^{T}\frac{\partial Loss}{\partial o^{i}_{t,n}}\frac{\partial o^{i}_{t,n}}{\partial u_{\text{th},n}},$ (20)
$\frac{\partial Loss}{\partial\tau_{n}}=\sum_{t=1}^{T}\frac{\partial Loss}{\partial u^{i}_{t,n}}\frac{\partial u^{i}_{t,n}}{\partial\tau_{n}}.$ (21)